PLEASE SHOW ALL STEPS

Olga owns a(n) race track that is worth 63,704 dollars and is expected to make annual cash flows forever. The cost of capital for the race track is 12.26 percent. The next annual cash flow is expected in 1 year and is expected to be 5,950 dollars. All subsequent cash flows are expected to grow annually at a constant growth rate. What is the cash flow produced by the race track in 6 years expected to be?

What is the value of an investment that will pay investors 4,200 dollars per year for 9 years and will also pay investors an additional 7,480 dollars in 1 year(s) from today if the expected return for the investment is 5.98 percent per year and the first annual payment of 4,200 dollars will be paid to investors in one year from today?
Answer #1

1.
Growth rate=Cost of capital-Next Annual cash flow/Present Value=12.26%-5950/63704=2.9199%
Cash flow in 6 years=(1+growth rate)^5*Next cash flow=(1+2.9199%)^5*5950=$6,870.90

2.
Value of the investment=4200/5.98%*(1-1/1.0598^9)+7480/1.0598=$35,649.93